# Adaptive Parameter-Free Robust Learning using Latent Bernoulli Variables

### v1

@@ 4.3 Overparameterized model
Dataset Method Symmetric Asymmetric Pairflip
20% 45% 20% 45% 20% 45%
Standard 95.66$\pm$0.28 87.47$\pm$0.82 98.29$\pm$0.14 87.73$\pm$0.65 97.21$\pm$0.41 69.26$\pm$3.65
Co-teaching 96.62$\pm$0.08 96.68$\pm$0.07 95.94$\pm$0.06 93.40$\pm$0.42 96.19$\pm$0.10 92.91$\pm$0.43
MNIST JoCoR $\bm{\pm}$0.05 98.38$\pm$0.10 99.02$\pm$0.05 95.21$\pm$3.70 98.96$\pm$0.09 91.06$\pm$5.01
CDR 98.81$\pm$0.09 98.27$\pm$0.09 $\bm{\pm}$0.05 94.35$\pm$1.29 98.95$\pm$0.06 87.55$\pm$1.26
USDNL 98.38$\pm$0.12 97.72$\pm$0.13 98.32$\pm$0.12 95.83$\pm$0.64 98.23$\pm$0.09 89.94$\pm$1.62
RLVI $\bm{\pm}$0.06 $\bm{\pm}$0.18 98.90$\pm$0.22 $\bm{\pm}$0.05 $\bm{\pm}$0.07 $\bm{\pm}$0.07
Standard 83.36$\pm$0.35 60.22$\pm$0.28 87.26$\pm$0.40 75.03$\pm$0.28 81.14$\pm$0.43 52.30$\pm$0.85
Co-teaching 87.55$\pm$0.38 83.70$\pm$0.46 86.98$\pm$0.26 64.84$\pm$0.87 86.57$\pm$0.21 67.38$\pm$2.93
CIFAR10 JoCoR 90.42$\pm$0.07 86.33$\pm$0.23 90.60$\pm$0.19 $\bm{\pm}$2.56 89.25$\pm$0.30 72.53$\pm$3.57
CDR 85.57$\pm$0.38 77.83$\pm$0.30 87.98$\pm$0.61 75.99$\pm$1.95 87.34$\pm$0.34 68.13$\pm$2.03
USDNL 88.65$\pm$0.21 83.33$\pm$0.23 88.36$\pm$0.33 74.62$\pm$0.70 87.05$\pm$0.24 66.49$\pm$2.45
RLVI $\bm{\pm}$0.14 $\bm{\pm}$0.33 $\bm{\pm}$0.12 $\bm{\pm}$0.50 $\bm{\pm}$0.27 $\bm{\pm}$0.29
Standard 61.96$\pm$0.11 42.36$\pm$0.80 62.88$\pm$0.20 39.91$\pm$0.58 62.62$\pm$0.53 38.94$\pm$0.42
Co-teaching 58.74$\pm$0.66 48.11$\pm$1.01 54.32$\pm$0.22 34.53$\pm$0.68 56.27$\pm$0.60 34.47$\pm$0.97
CIFAR100 JoCoR $\bm{\pm}$0.25 62.37$\pm$0.94 62.93$\pm$0.60 38.68$\pm$0.83 65.90$\pm$0.35 40.94$\pm$0.70
CDR 61.57$\pm$0.41 46.31$\pm$0.81 62.89$\pm$0.30 39.47$\pm$0.75 61.27$\pm$2.23 38.55$\pm$0.22
USDNL 64.96$\pm$0.56 53.82$\pm$1.15 59.12$\pm$0.10 36.44$\pm$0.93 61.76$\pm$1.01 35.81$\pm$0.20
RLVI $\bm{\pm}$0.55 $\bm{\pm}$0.79 $\bm{\pm}$0.83 $\bm{\pm}$2.12 $\bm{\pm}$0.77 $\bm{\pm}$1.01
@@
The results are presented as mean $\pm$ standard deviation over five runs with random initialization of network’s parameters. Table 2 shows that RLVI, with the described regularization, attains results competitive with alternative approaches. Also, because the number of training samples $n$ in deep learning is typically much smaller than the number of network’s parameters in $\bm{\theta}$, the additional steps in Algorithm 2 that involve $\bm{\pi}$ (fixed-point iterations and truncation) do not significantly increase the computational time compared to standard SGD. Table 3 shows the average time per epoch when training on CIFAR100.
@@
Table 3: Time per one epoch (in seconds) when training on corrupted CIFAR100: mean $\pm$ standard deviation over 200 epochs.
Standard Co-teaching JoCoR
8.82$\pm$0.08 17.91$\pm$0.35 18.11$\pm$0.11
CDR USDNL RLVI
13.88$\pm$0.35 9.40$\pm$0.22 10.52$\pm$0.56
@@
Real noise. To demonstrate how RLVI performs in a naturally noisy setting, we train ResNet50 (pre-trained on ImageNet [21]) on the challenging dataset Food101 [4] consisting of 101 food categories. For each class, 250 testing images were cleaned manually, while the remaining 750 training images per class still contain outliers. We use the Adam optimizer, a multi-step learning rate schedule (starting from $10^{-3}$), weight decay $10^{-4}$, and batch-size 32.
@@
In the real setting, the true $\varepsilon$ is unknown and has to be optimized for Co-teaching, JoCoR, CDR, and USDNL. Figure 6 shows that by varying $\varepsilon$ for alternative methods (from a large 20% – to a moderate 3% value), one can increase model performance. In contrast, RLVI improves over the standard approach without the need for additional optimization of the hyperparameter. In the Appendix, we provide results for other values of $\varepsilon$ for comparison.
Our implementation of the described experiments, including existing methods used for comparison, is available at github.com/akarakulev/rlvi.
@@
5 Conclusions
We presented the novel robust learning algorithm RLVI for likelihood maximization problems with corrupted datasets. It leverages variational inference to identify corrupted samples using latent Bernoulli variables. This alleviates the need for specifying hyperparameters such as the corruption level, which existing approaches rely on. RLVI can also be implemented as stochastic optimization, which makes it adaptive and applicable to learning from data with varying noise and out-of-core inference for large datasets. We demonstrated the effectiveness of the method on benchmark test problems in both – traditional statistical learning, as well as online and deep learning settings. The proposed RLVI algorithm meets or exceeds performance across considered experimental settings in a parameter-free and efficient manner.
Acknowledgements
The computations/data handling were enabled by the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre, and by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at Chalmers e-Commons at Chalmers, and Uppsala Multidisciplinary Center for Advanced Computational Science (UPPMAX) at Uppsala University, partially funded by the Swedish Research Council through grant agreement nos. 2022-06725 and 2018-05973. DZ acknowledges support from the Swedish Research Council through grant agreement no. 2018-05040.
Supplementary Material
Optimization w.r.t. probabilities $\bm{\pi}$ is convex
[…]
$\nabla^{2}_{\bm{\pi}}\mathcal{L}=\text{diag}\left(\dfrac{1}{\bm{\pi}(1-\bm{\pi})}\right)-\dfrac{\bm{1}\cdot\bm{1}^{\top}}{n\langle\bm{\pi}\rangle\langle 1-\bm{\pi}\rangle}.$ (21)
Convexity in $\bm{\pi}$ requires $\nabla^{2}_{\bm{\pi}}\mathcal{L}$ to be positive semi-definite, i.e., for any $\bm{x}\neq 0$,
$\bm{x}^{\top}\nabla^{2}_{\bm{\pi}}\mathcal{L}\,\bm{x}=\sum\limits_{i=1}^{n}x_{i}^{2}\cdot\dfrac{n\langle\bm{\pi}\rangle\langle 1-\bm{\pi}\rangle-\pi_{i}(1-\pi_{i})}{\pi_{i}(1-\pi_{i})n\langle\bm{\pi}\rangle\langle 1-\bm{\pi}\rangle}\geq 0.$ (22)
Given that $\bm{\pi}\in[0;1]^{n}$, the denominator cannot be negative, and therefore we only require to verify the sign of
$\sum\limits_{i=1}^{n}\Big{[}n\langle\bm{\pi}\rangle\langle 1-\bm{\pi}\rangle-\pi_{i}(1-\pi_{i})\Big{]}$ (23)
$=\sum\limits_{i=1}^{n}\pi_{i}\sum\limits_{j=1}^{n}(1-\pi_{j})-\sum\limits_{i=1}^{n}\pi_{i}+\sum\limits_{i=1}^{n}\pi_{i}^{2}$ (24)
$=(n-1)\sum\limits_{i=1}^{n}\pi_{i}-\sum_{i,j=1}^{n}\pi_{i}\pi_{j}+\sum\limits_{i=1}^{n}\pi_{i}^{2}$ (25)
$=\sum\limits_{i=1}^{n}\pi_{i}\Big{[}n-1-\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\pi_{j}\Big{]}\stackrel{{\scriptstyle?}}{{\geq}}0.$ (28)
Here, ‘?’ sign implies that we need to check if the final expression is not negative to prove positive semi-definiteness. However, in the last expression, $\pi_{i}\geq 0$ and $n-1\geq\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\pi_{j}$ for any $i=1,...,n$. Therefore, $\nabla^{2}_{\bm{\pi}}\mathcal{L}$ is positive semi-definite, and minimization of $\mathcal{L(\bm{\theta},\bm{\pi})}$ w.r.t. $\bm{\pi}$ is a convex optimization problem.
Convexity of $\mathcal{L}(\bm{\theta},\bm{\pi})$ in $\bm{\pi}$ allows for efficient computation of Bernoulli probabilities $\pi_{i},\,i=1,\dots,n$. In our implementation, we use the fixed-point algorithm to find the solution for the corresponding first-order necessary conditions $\nabla_{\bm{\pi}}\mathcal{L}=0$. Another option would be, for example, to use the Newton’s method, which requires inverting the Hessian $\nabla^{2}_{\bm{\pi}}\mathcal{L}$. But this matrix has rank one, and therefore can be inverted efficiently in $O(n)$ operations using the Woodbury formula. Also note that, when using the fixed-point algorithm (17), one needs to compute the exponential of the loss value for each sample. To avoid numerical issues and ensure convergence, we scale the range of the loss values with a constant (we divide by the average loss).
Online learning with varying level of noise
@@ Acknowledgements
In Figure 7, we provide the performance metrics of the standard SGD approach, RRM, and RLVI, used for binary classification of data that arrives in batches with varying number of outliers. In this figure, we show both accuracy and recall values for all three methods.
Figure 7: Online classification. Accuracy (left) and recall (right) on left-out data versus total number of observed samples – the metric is smoothed with a moving average filter.
Image classification with synthetic noise

### v2

@@ 4.3 Overparameterized model
Dataset Method Symmetric Asymmetric Pairflip Instance
20% 45% 20% 45% 20% 45% 20% 45%
Standard 95.66$\pm$0.28 87.47$\pm$0.82 98.29$\pm$0.14 87.73$\pm$0.65 97.21$\pm$0.41 69.26$\pm$3.65 95.66$\pm$0.36 71.98$\pm$1.43
Co-teaching 96.62$\pm$0.08 96.68$\pm$0.07 95.94$\pm$0.06 93.40$\pm$0.42 96.19$\pm$0.10 92.91$\pm$0.43 96.49$\pm$0.23 95.39$\pm$0.59
MNIST JoCoR 99.07$\pm$0.05 98.38$\pm$0.10 99.02$\pm$0.05 95.21$\pm$3.70 98.96$\pm$0.09 91.06$\pm$5.01 99.03$\pm$0.05 97.81$\pm$0.35
CDR 98.81$\pm$0.09 98.27$\pm$0.09 99.14$\pm$0.05 94.35$\pm$1.29 98.95$\pm$0.06 87.55$\pm$1.26 98.25$\pm$0.16 88.27$\pm$1.92
USDNL 98.38$\pm$0.12 97.72$\pm$0.13 98.32$\pm$0.12 95.83$\pm$0.64 98.23$\pm$0.09 89.94$\pm$1.62 98.13$\pm$0.11 96.52$\pm$0.38
BARE 99.07$\pm$0.11 $\bm{\pm}$0.10 $\bm{\pm}$0.06 $\bm{\pm}$0.12 99.05$\pm$0.06 98.22$\pm$0.21 99.05$\pm$0.07 $\bm{\pm}$0.45
RLVI $\bm{\pm}$0.06 $\bm{\pm}$0.18 98.90$\pm$0.22 $\bm{\pm}$0.05 $\bm{\pm}$0.07 $\bm{\pm}$0.07 $\bm{\pm}$0.03 98.38$\pm$0.08
Standard 83.36$\pm$0.35 60.22$\pm$0.28 87.26$\pm$0.40 75.03$\pm$0.28 81.14$\pm$0.43 52.30$\pm$0.85 81.76$\pm$0.37 55.51$\pm$1.04
Co-teaching 87.55$\pm$0.38 83.70$\pm$0.46 86.98$\pm$0.26 64.84$\pm$0.87 86.57$\pm$0.21 67.38$\pm$2.93 86.40$\pm$0.29 66.39$\pm$7.16
CIFAR10 JoCoR 90.42$\pm$0.07 86.33$\pm$0.23 90.60$\pm$0.19 76.94$\pm$2.56 89.25$\pm$0.30 72.53$\pm$3.57 89.08$\pm$0.36 79.48$\pm$1.37
CDR 85.57$\pm$0.38 77.83$\pm$0.30 87.98$\pm$0.61 75.99$\pm$1.95 87.34$\pm$0.34 68.13$\pm$2.03 85.72$\pm$0.74 66.82$\pm$2.48
USDNL 88.65$\pm$0.21 83.33$\pm$0.23 88.36$\pm$0.33 74.62$\pm$0.70 87.05$\pm$0.24 66.49$\pm$2.45 86.90$\pm$0.18 71.04$\pm$4.17
BARE 85.67$\pm$0.46 69.90$\pm$2.06 87.38$\pm$0.24 $\bm{\pm}$1.10 84.84$\pm$0.77 57.18$\pm$2.81 85.08$\pm$0.73 69.00$\pm$2.05
RLVI $\bm{\pm}$0.14 $\bm{\pm}$0.33 $\bm{\pm}$0.12 76.96$\pm$0.50 $\bm{\pm}$0.27 $\bm{\pm}$0.29 $\bm{\pm}$0.33 $\bm{\pm}$1.57
Standard 61.96$\pm$0.11 42.36$\pm$0.80 62.88$\pm$0.20 39.91$\pm$0.58 62.62$\pm$0.53 38.94$\pm$0.42 62.87$\pm$0.32 43.17$\pm$0.63
Co-teaching 58.74$\pm$0.66 48.11$\pm$1.01 54.32$\pm$0.22 34.53$\pm$0.68 56.27$\pm$0.60 34.47$\pm$0.97 57.44$\pm$0.32 34.97$\pm$0.61
CIFAR100 JoCoR $\bm{\pm}$0.25 62.37$\pm$0.94 62.93$\pm$0.60 38.68$\pm$0.83 65.90$\pm$0.35 40.94$\pm$0.70 67.98$\pm$0.32 52.47$\pm$0.31
CDR 61.57$\pm$0.41 46.31$\pm$0.81 62.89$\pm$0.30 39.47$\pm$0.75 61.27$\pm$2.23 38.55$\pm$0.22 62.09$\pm$0.45 41.80$\pm$0.77
USDNL 64.96$\pm$0.56 53.82$\pm$1.15 59.12$\pm$0.10 36.44$\pm$0.93 61.76$\pm$1.01 35.81$\pm$0.20 63.13$\pm$0.54 45.42$\pm$1.30
BARE 59.59$\pm$1.12 46.56$\pm$1.10 52.91$\pm$1.28 29.48$\pm$1.12 53.29$\pm$1.57 30.27$\pm$1.14 56.47$\pm$0.84 37.24$\pm$1.27
RLVI $\bm{\pm}$0.55 $\bm{\pm}$0.79 $\bm{\pm}$0.83 $\bm{\pm}$2.12 $\bm{\pm}$0.77 $\bm{\pm}$1.01 $\bm{\pm}$0.84 $\bm{\pm}$1.29
Standard 59.41$\pm$0.40 37.84$\pm$0.79 61.01$\pm$0.37 39.13$\pm$0.26 60.88$\pm$0.33 38.88$\pm$0.52
Co-teaching 59.77$\pm$0.67 48.45$\pm$1.44 55.64$\pm$0.90 35.98$\pm$0.82 58.44$\pm$1.27 35.61$\pm$0.47
CIFAR80N-O JoCoR 70.02$\pm$0.80 62.29$\pm$0.41 64.12$\pm$0.19 40.13$\pm$0.40 67.11$\pm$0.83 41.85$\pm$0.29
CDR 56.08$\pm$0.98 44.46$\pm$1.03 58.16$\pm$1.47 36.76$\pm$1.28 57.96$\pm$0.57 36.16$\pm$1.41
USDNL 64.07$\pm$1.63 52.00$\pm$2.89 59.43$\pm$0.69 37.04$\pm$0.46 62.20$\pm$0.52 37.12$\pm$1.09
BARE 57.39$\pm$1.23 42.56$\pm$2.20 54.47$\pm$1.01 30.12$\pm$2.09 55.21$\pm$1.28 30.09$\pm$0.86
RLVI $\textbf{71.13}\bm{\pm}\textbf{0.71}$ $\textbf{63.18}\bm{\pm}\textbf{0.36}$ $\textbf{71.96}\bm{\pm}\textbf{0.39}$ $\textbf{54.49}\bm{\pm}\textbf{1.76}$ $\textbf{71.45}\bm{\pm}\textbf{0.33}$ $\textbf{56.12}\bm{\pm}\textbf{0.23}$
@@
Results are presented as mean $\pm$ standard deviation over five runs with random initialization of network’s parameters. Table 2 shows that RLVI, with the described regularization, attains results competitive with alternative approaches. Also note that additional steps in Algorithm 2 involving $\bm{\pi}$ (fixed-point iterations and truncation) do not significantly increase the computational time compared to standard SGD. Table 3 shows the average time per epoch during training.
@@
Table 3: Time per one epoch in seconds when training on CIFAR100: mean $\pm$ standard deviation over 200 epochs. (Standard: 8.82$\pm$0.08)
Co-teaching JoCoR CDR
17.91$\pm$0.35 18.11$\pm$0.11 13.88$\pm$0.35
USDNL BARE RLVI
9.40$\pm$0.22 9.94 $\pm$ 0.10 10.52$\pm$0.56
@@
Real corruption. To demonstrate how RLVI performs in a naturally contaminated setting, we train ResNet50 – pre-trained on ImageNet (Wightman, 2019) – on the challenging dataset Food101 (Bossard et al., 2014) consisting of 101 food categories. For each class, 250 testing images were cleaned manually, while the remaining 750 training images per class still contain corrupted labels. We use the Adam optimizer with hyperparameters listed in Table 4.
@@
In the real setting, $\varepsilon$ is unknown and has to be optimized for Co-teaching, JoCoR, CDR, and USDNL. Figure 6 shows that by varying the estimated $\varepsilon$ for alternative methods (from a large 20% – to a moderate 3% value), one can enhance performance. In contrast, RLVI improves over the standard approach without the need for additional optimization of the hyperparameter. In Appendix, we provide results for other assumptions on $\varepsilon$.
Figure 6: Food101. Test accuracy assuming high and low $\varepsilon$.
5 Conclusions
We presented the novel robust learning algorithm RLVI for likelihood maximization problems with corrupted datasets. It leverages variational inference to identify corrupted samples under the Huber contamination model using latent Bernoulli variables. This alleviates the need for specifying hyperparameters such as the corruption level, which existing approaches rely on. RLVI can also be implemented as stochastic optimization, which makes it adaptive and applicable to learning from data with varying noise and out-of-core inference for large datasets. We demonstrated the effectiveness of the method on benchmark test problems in both – traditional statistical learning, as well as online and deep learning settings. The proposed RLVI algorithm meets or exceeds performance across considered experimental settings in a parameter-free and efficient manner.
Acknowledgements
The authors are thankful to Prof. Peter Stoica for insightful discussions during the work on this article. The computations/data handling were enabled by the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre and by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at Chalmers e-Commons at Chalmers, and Uppsala Multidisciplinary Center for Advanced Computational Science (UPPMAX) at Uppsala University, partially funded by the Swedish Research Council through grant agreement nos. 2022-06725 and 2018-05973. DZ and PS acknowledge support from the Swedish Research Council through grant agreement nos. 2018-05040 and 2023-05593 respectively.
Impact Statement
We introduce a novel approach for robust machine learning in likelihood maximization settings. The approach will potentially make robust machine learning easier to apply, owing to its parameter-free and general nature. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
A. Optimization of Bernoulli probabilities
A1. Convexity
Objective $\mathcal{L}(\bm{\theta},\bm{\pi})$ defined in Equation 15 is a convex function of variables $\bm{\pi}\in(0;1)^{n}$.
Proof.
The objective can be viewed as a sum of three terms,
$\mathcal{L}(\bm{\theta},\bm{\pi})=\sum\limits_{i=1}^{n}\pi_{i}\ell_{\bm{\theta}}(\bm{z}_{i})+\sum\limits_{i=1}^{n}\pi_{i}\ln\frac{\pi_{i}}{\langle\bm{\pi}\rangle}+\sum\limits_{i=1}^{n}(1-\pi_{i})\ln\frac{1-\pi_{i}}{1-\langle\bm{\pi}\rangle}.$ (23)
The first term is linear in $\bm{\pi}$, and we only need to verify if the second and the third terms are convex. Therefore, we first focus on the second term and its Hessian. Using $\bm{1}$ for an $n$-dimensional vector of ones and $\textbf{diag}(\cdot)$ for a diagonal matrix, we write
$f(\bm{\pi})$ $:=\sum\limits_{i=1}^{n}\pi_{i}\ln{\dfrac{\pi_{i}}{\langle\bm{\pi}\rangle}},$ (24)
$\nabla^{2}f(\bm{\pi})$ $=\textbf{diag}\left(\dfrac{1}{\pi_{i}}\right)-\dfrac{\bm{1}\cdot\bm{1}^{\top}}{n\langle\bm{\pi}\rangle}.$ (25)
For $f(\bm{\pi})$ to be convex, its Hessian has to be positive semi-definite: $\nabla^{2}f\succcurlyeq 0$. For clarity we multiply $\nabla^{2}f$ by a positive scalar $n\langle\bm{\pi}\rangle$ and consider a matrix
$n\langle\bm{\pi}\rangle\nabla^{2}f=\textbf{diag}\left(\dfrac{n\langle\bm{\pi}\rangle}{\pi_{i}}\right)-\bm{1}\cdot\bm{1}^{\top}=\bm{D}-\bm{1}\cdot\bm{1}^{\top}.$ (26)
Notice that $\bm{D}-\bm{1}\cdot\bm{1}^{\top}$ is a Schur complement of the block matrix
$\begin{pmatrix}\bm{D}&\bm{1}\\ \bm{1}^{\top}&1\end{pmatrix}.$ (27)
And, since $\bm{D}\succ 0$, the following inequalities should hold simultaneously by the properties of the Schur complement:
$\bm{D}-\bm{1}\cdot\bm{1}^{\top}\succcurlyeq 0$ $\iff$ $\begin{pmatrix}\bm{D}&\bm{1}\\ \bm{1}^{\top}&1\end{pmatrix}\succcurlyeq 0$ $\iff$ $1-\bm{1}^{\top}\bm{D}^{-1}\bm{1}\geq 0$ (28)
However, expression on the right holds true. Indeed,
$0\displaystyle 1-\bm{1}^{\top}\bm{D}^{-1}\bm{1}=1-\bm{1}^{\top}\textbf{diag}\left(\dfrac{\pi_{i}}{n\langle\bm{\pi}\rangle}\right)\bm{1}=1-\dfrac{1}{n\langle\bm{\pi}\rangle}\sum\limits_{i=1}^{n}\pi_{i}=1-\dfrac{\langle\bm{\pi}\rangle}{\langle\bm{\pi}\rangle}=$ 0 . (29)
Therefore, we conclude that $\nabla^{2}f\succcurlyeq 0$ and the second term in $\mathcal{L}(\bm{\theta},\bm{\pi})$ is convex. To establish the convexity of the third term, it suffices to consider $1-\pi_{i}$ as its variables, which makes the proof identical to the above. Hence $\mathcal{L}(\bm{\theta},\bm{\pi})$ is convex in $\bm{\pi}$ as a sum of convex functions. ∎
A2. Convergence
@@ A. Optimization of Bernoulli probabilities
B. Online learning with varying level of noise
@@
In Figure 7, we provide the performance metrics of the standard SGD approach, RRM, and RLVI, used for binary classification of data that arrives in batches with varying number of corrupted labels. In this figure, we show both accuracy and recall values for all three methods. The metrics are smoothed with a moving average filter.
To obtain the main results in online classification, shown in Figure 7, we used batches of 100 observations during optimization (and reserved 100 samples for computing accuracy and recall in each iteration). To demonstrate how batch size affects the performance, in Figures 8 and 9 we additionally show the same metrics when 75 and 50 examples are used for training respectively (but 100 samples are still used for testing).
Figure 7: Online classification. Accuracy and recall when learning with the batch size 100.
Figure 8: Online classification. Accuracy and recall when learning with the batch size 75.
Figure 9: Online classification. Accuracy and recall when learning with the batch size 50.
Table 4: Hyperparameter settings for the deep learning experiments (LR = learning rate, mom. = momentum)
Dataset MNIST CIFAR10 CIFAR100 & CIFAR80N-O Food101
Model LeNet ResNet18 ResNet34 ResNet50 (pre-trained on Imagenet)
Optimizer SGD with mom. 0.9 SGD with mom. 0.9 SGD with mom. 0.9 Adam
Epochs 100 200 200 80
Batch size 32 128 128 32
Weight decay $10^{-3}$ $5\cdot 10^{-4}$ $5\cdot 10^{-4}$ $10^{-4}$
LR schedule linear decay cosine annealing cosine annealing multi-step
Initial LR $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-3}$
C. Overparameterized setting: image classification using convolutional neural networks
C1. Hyperparameters
In the experiments on image classification, we use hyperparameter configurations that follow the common settings found in literature. These settings are listed in Table 4.
C2. Synthetic corruption